# HippoMM: Hippocampal-inspired Multimodal Memory for Long Audiovisual Event Understanding

### v1

@@ Abstract
Comprehending extended audiovisual experiences remains a fundamental challenge for computational systems. Current approaches struggle with temporal integration and cross-modal associations that humans accomplish effortlessly through hippocampal-cortical networks. We introduce HippoMM, a biologically-inspired architecture that transforms hippocampal mechanisms into computational advantages for multimodal understanding. HippoMM implements three key innovations: (i) hippocampus-inspired pattern separation and completion specifically designed for continuous audiovisual streams, (ii) short-to-long term memory consolidation that transforms perceptual details into semantic abstractions, and (iii) cross-modal associative retrieval pathways enabling modality-crossing queries. Unlike existing retrieval systems with static indexing schemes, HippoMM dynamically forms integrated episodic representations through adaptive temporal segmentation and dual-process memory encoding. Evaluations on our challenging HippoVlog benchmark demonstrate that HippoMM significantly outperforms state-of-the-art approaches (78.2% vs. 64.2% accuracy) while providing substantially faster response times (20.4s vs. 112.5s). Our results demonstrate that translating neuroscientific memory principles into computational architectures provides a promising foundation for next-generation multimodal understanding systems. The code and benchmark dataset are publicly available at https://github.com/linyueqian/HippoMM.
@@ 1 Introduction
Imagine hearing a familiar melody through an open window, instantly recalling not merely the music but the complete concert experience: the interplay of stage lights, the energy of the crowd, and even the aroma of summer rain. This remarkable cognitive capability, known as multimodal pattern completion, involves reconstructing integrated experiences from partial, cross-modal stimuli and remains a significant frontier in computational modeling. Although current methods proficiently process static multimodal data [radford2021learning, girdhar2023imagebind] or retrieve relevant segments from extensive video corpora [ren2024videorag], they struggle fundamentally with long-form multimodal comprehension. Specifically, these approaches exhibit limitations in three key dimensions: segmenting continuous audiovisual streams into coherent episodic units, reconstructing rich experiences by integrating cross-modal and temporal information (particularly from partial cues), and flexibly retrieving information across modalities—such as recalling visual details from auditory cues.
In contrast, the human hippocampus robustly achieves these capabilities. As illustrated conceptually in Figure 1 (top), the hippocampus integrates multisensory inputs (e.g., visual and auditory streams) through structures including the entorhinal cortex, organizes continuous experiences into discrete episodes [baldassano2017discovering], employs pattern separation mechanisms typically attributed to the Dentate Gyrus (DG), and performs pattern completion associated primarily with CA3 circuits to store and recall distinct memory traces [yassa2011pattern, rolls2013quantitative]. These processes ultimately yield cohesive episodic representations, mediated by hippocampal circuitry including the CA1 region [eichenbaum2014time].
Figure 1: Conceptual overview of hippocampal versus HippoMM multimodal processing. (Top) Biological hippocampus integrates visual and auditory modalities through the entorhinal cortex and associated circuits (DG, CA3, CA1) to form and recall episodic memories. (Bottom) Proposed HippoMM architecture processes multimodal inputs (video, audio, cross-modal) inspired by hippocampal principles for episodic memory formation and cue-driven retrieval.
Despite this clear biological blueprint, existing computational memory systems often mimic only isolated aspects. They typically operate unimodally [gutierrez2024hipporag, graves2016hybrid] or function as passive retrieval systems [lewis2020retrieval, ren2024videorag], falling short of replicating the active, reconstructive, and crossmodal associative nature of hippocampal memory, particularly for extended audiovisual data streams. To address these limitations, we introduce HippoMM, a novel architecture designed to computationally implement functional principles inspired by hippocampal mechanisms for long form multimodal content understanding, as conceptualized in Figure 1 (bottom). Processing video, audio, and jointly learned multimodal representations, HippoMM aims to emulate the brain’s ability to form and query rich, integrated multimodal memories. While prior work has drawn inspiration from the hippocampus for unimodal tasks or specific memory functions, HippoMM represents, to our knowledge, the first architecture to systematically adapt and integrate a suite of hippocampal memory mechanisms to address the unique challenges of long-form audiovisual understanding. Our contributions are as follows:
  1. 1.
    Hippocampal Dynamics for Long AV Streams: We operationalize dentate gyrus-inspired pattern separation and CA3-like pattern completion specifically for continuous audio-visual data. This allows HippoMM to segment lengthy streams into distinct episodes and reconstruct coherent multimodal experiences even from fragmented cues, moving beyond passive retrieval.
  2. 2.
    Short-to-Long Term Memory Consolidation: We model the transition from detailed perceptual traces (short-term) to abstract semantic representations (long-term) through a neuro-inspired consolidation and semantic replay process. This contrasts with models that use fixed chunking or lack mechanisms for evolving representations over time, enabling more efficient and meaningful long-term storage.
  3. 3.
    Cross-Modal Associative Querying: HippoMM introduces retrieval pathways that explicitly support querying information in one modality using cues from another (e.g., finding what was seen when a specific sound occurred, or vice-versa). This capability, absent in prior multimodal systems that often disregard audio richness or lack such cross-modal retrieval tests, mimics a key aspect of human associative recall. Unlike existing approaches that treat modalities in isolation, our system maintains the temporal co-occurrence relationships essential for reconstructing coherent multimodal experiences from partial cues.
Evaluated on our proposed HippoVlog benchmark, comprising long-form vlogs with challenging cross-modal and unimodal question-answering tasks, HippoMM achieves state-of-the-art accuracy (78.2% average accuracy, significantly outperforming VideoRAG [ren2024videorag]), particularly excelling in cross-modal and visual reasoning tasks. Ablation studies underscore the critical roles of biologically inspired components: removing detailed recall mechanisms reduces visual accuracy by 6%, and eliminating adaptive reasoning results in a 1.6% decline in semantic understanding. The remainder of the paper is structured as follows: Section 2 discusses related research on hippocampal mechanisms and multimodal understanding; Section 3 formalizes the multimodal memory-augmented retrieval task and introduces the HippoVlog benchmark; Section 4 describes the HippoMM architecture in detail; Section 5 presents comprehensive experimental results; and finally, we conclude with discussions on implications and future directions in Section 6. Our work demonstrates that integrating hippocampal principles can advance computational understanding of complex, continuous multimodal experiences, offering a robust foundation toward developing systems with human-like memory and reasoning capabilities.
@@ 2.1 Hippocampal Memory Mechanisms
Neuroscientific research highlights the hippocampus’s crucial ability to form, organize, and retrieve episodic memories. Foundational work established key principles including pattern separation through sparse coding in the dentate gyrus [yassa2011pattern], autoassociative pattern completion in CA3 [rolls2013quantitative], and theta-gamma oscillations for temporal coding [lisman2009prediction]. Recent findings reveal how the hippocampus segments experiences at event boundaries [baldassano2017discovering, ben2013hippocampal], creating hierarchical memory structures [zacks2007event] that optimize storage and recall. Computational models have translated these principles to machine learning through temporal chunking [dubrow2013influence] and neural memory networks [kumaran2016learning], but primarily in unimodal contexts. The hippocampus’s role as a multimodal integrator [teyler2007hippocampal, eichenbaum2014time] provides critical insights for artificial memory systems. Neurobiological studies demonstrate its capacity to bind cross-modal features while maintaining modality-specific details [quiroga2009explicit, mormann2008latency], a capability existing computational models lack. Critically, effectively integrating continuous, cross-modal audiovisual streams remains underdeveloped, even in recent memory architectures [whittington2020tolman], presenting an important research direction.
2.2 Audio-Visual Multimodal Understanding
Modern multimodal systems leverage joint embedding spaces [radford2021learning, girdhar2023imagebind] to align audio-visual modalities, while video-language models [alayrac2022flamingo, zhang2023video] employ temporal attention for sequential processing. Though effective for static retrieval, these methods lack mechanisms for temporal memory organization critical for long-form event understanding. Recent audio-visual models [shu2023audio, guo2025aligned] enhance modality integration through specialized tokens or multi-scale adapters, yet struggle with extended temporal contexts and cross-modal associative memory. Retrieval-augmented approaches [ren2024videorag, yuan2025memory] address long-context challenges through graph-based grounding and cognitive memory steps, but remain biologically uninspired. Cross-modal retrieval frameworks [lin2024mm] mitigate modality bias yet lack temporal organization, while efficiency-focused methods [lin2025speechprune, liu2025keyframe] prune redundant elements without enabling progressive retrieval pathways.
@@ 2.3 Biomimetic Memory Architectures
Memory-augmented neural networks have evolved from early differentiable memory systems [santoro2016meta] to sophisticated architectures explicitly mirroring hippocampal-cortical interactions [whittington2020tolman]. Hippocampus-inspired models like HippoRAG [gutierrez2024hipporag] adapt principles of hippocampal indexing theory for knowledge integration, drawing on biological mechanisms similar to the dentate gyrus’ pattern separation [yassa2011pattern] and CA3’s autoassociative recall [rolls2013quantitative] for efficient document retrieval. Similarly, Differentiable Neural Computers [graves2016hybrid] emulate the hippocampus’ sparse temporal coding mechanisms [lisman2009prediction]. However, these approaches remain predominantly unimodal, neglecting the hippocampus’ critical role as a multimodal binding engine that coordinates cortical sensory streams through theta-phase coding, grid cell-inspired spatial organization, and cross-modal integration [eichenbaum2014time].
@@ 3.1 Task Formulation
We formalize the Multimodal Memory-augmented Retrieval (MMR) task as the challenge of developing systems that process, understand, and reason about long-form continuous multimodal data streams $X=\{x_{t}\}$, exemplified by extended audiovisual recordings. Given a user query $q$ targeting visual ($q_{V}$), auditory ($q_{A}$), cross-modal ($q_{V\times A}$), or semantic ($q_{S}$) information, the system must construct and utilize appropriate memory representations to synthesize a coherent and accurate answer $a$. The task decomposes into three fundamental stages. First, a memory formation function $\mathcal{F}_{\text{mem}}$ transforms the raw input stream $X$ into a structured memory representation $M$:
$M=\mathcal{F}_{\text{mem}}(X)$ (1)
This memory $M=\{m_{k}\}$ comprises encoded segments $m_{k}$, each capturing multimodal features, temporal context, and semantic abstractions. Second, a retrieval function $\mathcal{R}_{\text{retrieval}}$ processes the query $q$ against memory $M$ to identify relevant evidence $E$:
$E=\mathcal{R}_{\text{retrieval}}(q,M)$ (2)
Here, $E$ typically includes a relevance-ranked subset of memory elements (i.e., $E=\{\langle m_{k},r_{k}\rangle\mid m_{k}\in M\land\text{relevant}(m_{k},q)\}$). Finally, an answer synthesis function $\mathcal{G}_{\text{synth}}$ generates the response $a$ by reasoning over the query $q$ and retrieved evidence $E$:
$a=\mathcal{G}_{\text{synth}}(q,E)$ (3)
Implementing an effective MMR system requires sophisticated realizations of $\mathcal{F}_{\text{mem}}$, $\mathcal{R}_{\text{retrieval}}$, and $\mathcal{G}_{\text{synth}}$ that handle temporal dependencies, enable cross-modal associations, and support robust reasoning, analogous to biological memory processes described in cognitive neuroscience [yassa2011pattern, eichenbaum2014time].
3.2 HippoVlog Dataset
To evaluate MMR systems under realistic conditions, we introduce HippoVlog. Existing benchmarks for multimodal understanding typically utilize short video clips, rely solely on Large Language Model (LLM)-based evaluation without ground truth answers [ren2024videorag], or lack audio elements for cross-modal reasoning within continuous, long-form videos [MLVU]. HippoVlog addresses these gaps, comprising 25 long-form daily vlogs (682 minutes total) with naturalistic audiovisual content. It includes 1,000 validated multiple-choice question-answer pairs, each with four candidate answers, targeting specific memory functions:
  • Cross-modal binding ($T_{V\times A}$): Linking visual/auditory cues (e.g., "What object appeared when the speaker mentioned hiking?").
  • Auditory-focused retrieval ($T_{A}$): Extracting audio details (e.g., "What instrument played during the introduction?").
  • Visual-focused retrieval ($T_{V}$): Extracting visual details (e.g., "What text was visible on the store sign?").
  • Semantic reasoning ($T_{S}$): Integrating semantic information over time (e.g., "What are the cooking steps mentioned throughout the video?").
Questions were created using GPT-4o [hurst2024gpt], manually validated, and annotated with timestamps. HippoVlog’s objective assessment format, focus on audio-visual integration, and long-form temporal dynamics make it particularly suitable for evaluating MMR systems. Construction details are provided in the Appendix and our dataset is publicly available [^1] https://github.com/linyueqian/HippoVlog/.
Figure 2: The HippoMM architecture for multimodal memory. (a) Memory Formation: composed of Temporal Pattern Separation ($\mathcal{S}_{t}$) based on perceptual boundaries, Perceptual Encoding of visual and auditory inputs with cross-modal features, Memory Consolidation using similarity-based filtering ($K$), and Semantic Replay generating ThetaEvent representations ($\theta$). (b) Memory Retrieval: implementing Query-Driven Pattern Completion through fast retrieval ($\Phi_{\text{fast}}$) and detailed recall pathways with temporal window localization ($\mathbf{W}$).
4 HippoMM System
4.1 Overview
We introduce HippoMM, a computational architecture designed for the complex task of understanding and reasoning about long-form multimodal data streams, typified by extended audiovisual recordings. The architecture draws functional inspiration from the mammalian hippocampus’s established roles in episodic memory processing [eichenbaum2017prefrontal], including cross-modal binding, temporal sequence organization, pattern separation (differentiating similar inputs), pattern completion (retrieving full memories from partial cues), and memory consolidation. While HippoMM does not model neural dynamics directly, it seeks to replicate key functional principles through distinct algorithmic modules. Our objective is to overcome the limitations of conventional multimodal systems, which often lack robust mechanisms for handling continuous temporal dynamics and associative cross-modal recall. HippoMM enables context-sensitive retrieval, facilitating the reconstruction of integrated multimodal experiences from potentially incomplete queries (e.g., retrieving visual context associated with a specific sound event). The system architecture, depicted schematically in Figure 2, comprises two primary phases: Memory Formation (Sec. 4.2), which processes raw inputs into structured, abstracted representations in long-term memory (Figure 2a), and Memory Retrieval (Sec. 4.3), which leverages these representations to answer user queries via hierarchical and associative search strategies (Figure 2b).

### v2

@@ Abstract
Comprehending extended audiovisual experiences remains challenging for computational systems, particularly temporal integration and cross-modal associations fundamental to human episodic memory. We introduce HippoMM, a computational cognitive architecture that maps hippocampal mechanisms to solve these challenges. Rather than relying on scaling or architectural sophistication, HippoMM implements three integrated components: (i) Episodic Segmentation detects audiovisual input changes to split videos into discrete episodes, mirroring dentate gyrus pattern separation; (ii) Memory Consolidation compresses episodes into summaries with key features preserved, analogous to hippocampal memory formation; and (iii) Hierarchical Memory Retrieval first searches semantic summaries, then escalates via temporal window expansion around seed segments for cross-modal queries, mimicking CA3 pattern completion. These components jointly create an integrated system exceeding the sum of its parts. On our HippoVlog benchmark testing associative memory, HippoMM achieves state-of-the-art 78.2% accuracy while operating 5x faster than retrieval-augmented baselines. Our results demonstrate that cognitive architectures provide blueprints for next-generation multimodal understanding.
@@ 1 Introduction
Despite recent progress, current multimodal models still cannot perform long-form audiovisual understanding accurately and efficiently. Specifically, they (i) must retain entire continuous content spanning several hours as the context [42], (ii) cannot reconstruct complete experiences when given partial sensory cues [17], and (iii) cannot form abstract memories that transfer across contexts [35]. Such failures prevent machines from seamlessly understanding real-world experiences the way humans naturally do: we effortlessly (i) segment raw, continuous content into meaningful episodes [40], (ii) recall entire episodes from a single sound or image [33], and (iii) extract lasting knowledge from fleeting perceptions [23].
Figure 1: Conceptual Overview of Hippocampal vs. HippoMM Processing. The biological hippocampus (top) and our HippoMM architecture (bottom) both integrate continuous sensory streams to form and recall episodic memories.
While prior works improve multimodal systems by scaling up or introducing sophisticated architectures [26, 9], recent memory-augmented approaches [14] show explicit memory mechanisms are crucial. We posit that current systems lack effective cognitive mechanisms for episodic memory. The biological hippocampus provides a cognitive architecture that solves these challenges through three core mechanisms: segmentation of continuous experience into discrete episodes, associative reconstruction from fragmentary inputs, and consolidation of perceptual details into semantic knowledge [38, 29]. Figure 1 illustrates the parallel between biological hippocampal processing and our computational approach. By translating these principles into algorithmic implementations, we design an accurate and efficient multimodal system.
To this end, in this work we propose HippoMM, which implements our insight through direct functional mapping of hippocampal computations. Following dentate gyrus (DG) pattern separation [38], our Episodic Segmentation employs content-adaptive boundaries to parse inputs into episodic codes. Emulating hippocampal consolidation processes [16], Memory Consolidation progressively transforms perceptual traces into semantic representations, culminating in ThetaEvent memories. Mimicking CA3 autoassociation [29], Hierarchical Memory Retrieval enables pattern completion from partial cues and cross-modal queries. This integrated architecture amplifies each component through their interactions, achieving performance gains that validate the cognitive architectural approach.
Another contribution of our work is HippoVlog, a benchmark designed to test cross-modal associative recall in long-form video understanding. Unlike existing benchmarks that focus on sequential comprehension [22] or simple retrieval tasks [39], HippoVlog evaluates episodic memory functions through partial cue reconstruction and semantic consolidation questions. Built from 25 daily vlogs, it provides 1,000 manually-validated questions across four memory function categories.
In summary, our contributions are:
  • We design HippoMM, which implements hippocampal memory functions through episodic segmentation at perceptual boundaries, cross-modal feature binding in consolidated memory units, and hierarchical retrieval that balances semantic summaries with detailed recall.
  • We introduce HippoVlog, a benchmark for cross-modal associative memory with questions requiring pattern completion from partial cues, addressing the gap in existing benchmarks that only test sequential comprehension.
  • We demonstrate that cognitive architectures provide principled solutions: HippoMM achieves 78.2% accuracy and speedup over retrieval baselines, validating biological mechanisms as effective computational designs.
@@ 2.1 Hippocampal Memory Mechanisms
The hippocampus performs three key functions: pattern separation through sparse coding in the DG [38], pattern completion via CA3 autoassociative networks [29], and event segmentation at boundaries [3]. While computational models have adapted these principles [16], they primarily focus on unimodal contexts, overlooking the hippocampus’s role as a multimodal integrator [6].
2.2 AudioVisual Multimodal Understanding
Current multimodal systems employ joint embeddings [26, 9] or video-language models [1, 20] but lack mechanisms for temporal memory organization and associative recall. While recent models like VideoLLaMA 2 [5] and Qwen2.5-Omni [37] incorporate audiovisual processing, and voice-native benchmarks [19] expose modality-induced reasoning gaps, these efforts focus on modality alignment rather than memory formation. Related benchmarks such as LongVALE [8] target temporal grounding and dense captioning—predicting temporal boundaries in pre-segmented video—which is complementary to but distinct from our task of cross-modal associative recall from continuous streams. Critically, existing approaches cannot form episodic memories from continuous streams or perform cross-modal pattern completion, where sounds trigger visual memories or vice versa. This fundamental gap in hippocampal-like memory mechanisms represents a major limitation, as without them, models process only pre-segmented clips through attention [32] or fusion [24, 18, 34] without true associative recall.
@@ 2.3 Biomimetic Memory Architectures
Memory-augmented networks like the Differentiable Neural Computer [12] pioneered external memory but remain predominantly unimodal. Recent hippocampal-inspired architectures include HippoRAG [13] for text retrieval and MA-LMM [14] which extends memory-augmented approaches to long-term video understanding through memory banks. HippoMM uniquely integrates pattern separation, consolidation, and cross-modal pattern completion in a unified architecture for continuous audiovisual understanding.
Figure 2: The HippoMM architecture. (a) Memory Formation: composed of Episodic Segmentation ($\mathcal{S}_{t}$) based on perceptual boundaries, Perceptual Encoding of visual and auditory inputs with cross-modal features, and Memory Consolidation that progressively transforms short-term features through similarity-based filtering ($K$) into consolidated features, feature collections, and ultimately semantic summaries to form ThetaEvent ($\theta$). (b) Hierarchical Memory Retrieval: implementing Query-Driven Pattern Completion through fast retrieval ($\Phi_{\text{fast}}$) and detailed recall pathways with temporal window localization ($\mathbf{W}$).
@@ 3.1 Task Formulation
We formalize the task of Multimodal Memory-augmented Retrieval (MMR). While retrieval tasks have been extensively studied across various modalities, existing formulations typically treat retrieval as matching queries against pre-indexed content. Our work uniquely addresses retrieval from continuous multimodal streams requiring active memory formation, which is fundamentally different from traditional settings as the system must continuously construct and maintain structured memory representations from ongoing inputs, mimicking human episodic memory.
The objective is to develop a system that processes a long-form, continuous multimodal data stream $X=\{x_{t}\}$ and, given a query $q$, synthesizes an accurate answer $a$ through three functional stages: (i) memory formation $\mathcal{F}_{\text{mem}}$ transforms $X$ into structured representation $M$; (ii) retrieval $\mathcal{R}_{\text{retrieval}}$ identifies relevant evidence $E$ from $M$; and (iii) answer synthesis $\mathcal{G}_{\text{synth}}$ generates response $a$. HippoMM instantiates $\mathcal{F}_{\text{mem}}$ via content-adaptive segmentation, cross-modal encoding, and similarity-filtered consolidation; $\mathcal{R}_{\text{retrieval}}$ via GPT-4o-based confidence-gated hierarchical retrieval; and $\mathcal{G}_{\text{synth}}$ via LLM-based adaptive reasoning—each grounded in hippocampal principles but not the only valid instantiation.
3.2 The HippoVlog Benchmark
A significant challenge in evaluating MMR systems is the lack of benchmarks designed to test for cross-modal associative recall. Recent benchmarks have advanced long-form video understanding significantly: MLVU [43] evaluates multi-task comprehension across diverse genres, ALLVB [31] unifies nine tasks with 2-hour videos, and Video-MME [7] provides comprehensive multimodal evaluation. However, these benchmarks fundamentally test comprehension of presented information rather than memory formation and associative recall.
Specifically, existing benchmarks evaluate whether models can answer questions about explicitly shown content, but they do not test: (i) the ability to form episodic memories from continuous experience that enable later reconstruction from partial cues, (ii) cross-modal pattern completion where a sound triggers recall of associated visual content or vice versa, and (iii) the consolidation of perceptual details into transferable semantic knowledge. These capabilities, central to human episodic memory, require fundamentally different evaluation paradigms than sequential comprehension.
To address this gap, we introduce HippoVlog, a benchmark specifically designed to evaluate hippocampal-like memory functions. Built from 25 long-form daily vlogs totaling 682 minutes, it provides 1,000 questions across four memory function categories: Cross-modal binding ($T_{V\times A}$), Auditory-focused retrieval ($T_{A}$), Visual-focused retrieval ($T_{V}$), and Semantic reasoning ($T_{S}$). The quality of the benchmark is ensured through rigorous manual validation, achieving a high inter-annotator agreement (Cohen’s $\kappa=0.975$). Unlike comprehension-focused benchmarks, HippoVlog’s questions probe whether systems can reconstruct complete multimodal experiences from fragmentary cues—the hallmark of episodic memory. Construction details are provided in the Appendix. Our code is publicly available. [^1] Code: https://github.com/linyueqian/HippoMM
4 Methodology
We implement the fundamental cognitive architecture of the hippocampus, translating its three core computations of pattern separation (DG), pattern completion (CA3), and memory consolidation (CA1) into algorithms for multimodal understanding. This hippocampal-inspired architecture addresses precisely those challenges that confound current models: segmenting continuous experience into discrete episodes, reconstructing complete memories from partial cues, and abstracting semantic knowledge from perceptual details. Our methodology implements these biological principles through a two-phase process: a Memory Formation phase that mirrors the hippocampal encoding pathway from entorhinal cortex through DG-CA3-CA1, and a Hierarchical Memory Retrieval phase that implements CA3 autoassociation and CA1 semantic recall for query-driven reasoning.